OpenStudy (anonymous):

A tugboat goes 140 miles upstream in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.

OpenStudy (anonymous):

The tugboats speed and the speed of the current

OpenStudy (shane_b):

Start by knowing \[Speed=\frac{distance}{time}\]The speed upstream will be (where c = current):\[s-c=\frac{140mi}{10hr}=14mph\]The speed upstream will be (where c = current):\[s+c=\frac{140mi}{5hr}=28mph\]The difference between the two speeds will equal exactly half of the current's speed (since it helps you in one direction and hurts you in the other). Once you know that you should be able to clearly see what the boat's speed in still water will be.

OpenStudy (shane_b):

Note a typo: The second equation should have read "downstream"...I cut and pasted that and forgot to change the word.

OpenStudy (anonymous):

Thank you so much you made the equation sound so simple.

OpenStudy (shane_b):

yw :)

OpenStudy (anonymous):

The current is 21 and the speed is 7 that's the answer

OpenStudy (shane_b):

You have it backwards....and you're 8 months late :)